A cart is initially travelling across a flat surface with a velocity of 1.6 m/s. A net force of 87 N is 

exerted on the cart over a distance of 15.4 m for 4.1 s. What is the mass of the cart? 


1
Expert's answer
2021-05-11T13:59:41-0400

"s=v_0t+at^2/2\to a=\frac{2(s-v_0t)}{t^2}=\frac{2(15.4-1.6\cdot 4.1)}{4.1^2}=1.052\ (m/s^2)"


"F=ma\to m=F/a=87/1.052=82.7\ (kg)" . Answer